BexleyBrooks:

Solve for the letter. Express your answer as a proper or improper fraction in simplest terms. Problem #1: 1/12a-1/2=-1/8 what is a? Problem #2: -25+4/9b=-21 what is b? Rewrite in the simplest terms: −(−9q+3r)+6r−2(−4r+9q) What is the simplest term?

MikiJohnny:

........

MikiJohnny:

oml

MikiJohnny:

idek tbh

Bigrizzlertwizzler:

am kinda dumb 😰

MikiJohnny:

same

MikiJohnny:

so idrk

Bigrizzlertwizzler:

am cooked

MikiJohnny:

lol

idontexistrn1:

so don't solve it?

toga:

as promised @magaback

MikiJohnny:

idrk...u can try

idontexistrn1:

@toga wrote:
as promised @magaback
buuuuuuuut aren't u the calculator?

toga:

@idontexistrn1 wrote:
@toga wrote:
as promised @magaback
buuuuuuuut aren't u the calculator?
I am but made a promise to him that the next math I see I'll tag him cause I always answer them before he can

idontexistrn1:

ooooh ok

toga:

@magaback

MAGABACK:

@toga wrote:
@magaback
yes?

toga:

@toga wrote:
@idontexistrn1 wrote:
@toga wrote:
as promised @magaback
buuuuuuuut aren't u the calculator?
I am but made a promise to him that the next math I see I'll tag him cause I always answer them before he can
@magaback wrote:
@toga wrote:
@magaback
yes?

toga:

do you want to answer this one?

MAGABACK:

sure

MAGABACK:

ok I believe we need to do common denominators first for #1. So common denominator is 24. 2/24 -12/24 and -3/8 so the equation would be 2/24a-12/24=-3/24 then I think we add 12/24 to both sides to get 2/24a=9/24. Then I believe u divide by 2, and get a=4 1/24. right @toga?

MAGABACK:

I gtg so toga can help u with the others

BexleyBrooks:

@magaback wrote:
ok I believe we need to do common denominators first for #1. So common denominator is 24. 2/24 -12/24 and -3/8 so the equation would be 2/24a-12/24=-3/24 then I think we add 12/24 to both sides to get 2/24a=9/24. Then I believe u divide by 2, and get a=4 1/24. right @toga?
It was wrong or at least thats what it says

toga:

@magaback wrote:
ok I believe we need to do common denominators first for #1. So common denominator is 24. 2/24 -12/24 and -3/8 so the equation would be 2/24a-12/24=-3/24 then I think we add 12/24 to both sides to get 2/24a=9/24. Then I believe u divide by 2, and get a=4 1/24. right @toga?
It looks like you made a small mistake in your calculations. When you add 12/24 to both sides, you should get 2/24a = 9/24, which simplifies to 2/24a = 3/8, not 9/24. Then, when you divide by 2, you should get a = 3/4, not 4 1/24.